OpenStudy (anonymous):

a compass an a ruler together cost $2. the compass costs 90 cents more than the ruler how much does the compass cost

OpenStudy (anonymous):

c + r = 4.00 c = r + 90 so r + r + 90 = 4.00 2 * r = 3.10 r = 1.55 c = 1.55 + 0.90 = 2.45

OpenStudy (anonymous):

okay thank you